Circular No. 8907

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V459 VULPECULAE = NOVA VULPECULAE 2007 No. 2 S. Nakano, Sumoto, Japan, reports the discovery by Hiroshi Kaneda

(Sapporo, Japan) of an apparent nova (mag 8.7) on three 4-s unfiltered CCD frames taken on 2007 Dec. 25.35 UT using a Nikon D40 digital camera (+ 105-mm f/2.5 lens), with an apparent independent discovery made by A. Tago at mag 8.3 on his two patrol frames taken on Dec. 26.38 using a Canon 20D digital camera (+ Pentax 105-mm f/3.2 lens). Nakano’s measurement of the variable from Tago’s image yielded the position α = 19^{h}48^{m}08^{s}.81, δ = +21^{o}15′29′′.9 (equinox 2000.0) and mag 7.8. Additional CCD magnitudes: Dec. 8, [10–11 (Kaneda); 26.342, 7.1 (K. Itagaki, Yamagata, Japan, 0.60-m f/5.7 reflector; position end figures 08^{s}.84, 27′′.6); 26.355, 7.7 (Kaneda; position end figures 08^{s}.89, 26′′.8). Kaneda and K. Kadota (Ageo, Japan) note that a USNO-B1.0-catalogue star has position end figures 08^{s}.871, 26′′.81 and mag ∼ 20. Spectroscopy by numerous observers, as reported by H. Yamaoka and U. Munari on CBETs 1181 and 1183, confirm the object as a nova. Visual magnitude estimates for V459 Vul by A. Pereira, Cabo da Roca, Portugal: Dec. 26.79, 8.0; 27.79, 7.7; 30.80, 7.9.

V723 CASSIOPEIAE S. Mazuk, R. J. Rudy, D. K. Lynch, A. M. Gilbert, and T. R. Prater,

The Aerospace Corporation; R. B. Perry, Langley Research Center, NASA; R. C. Puetter, University of California, San Diego; G. Schwarz, West Chester University; and C. E. Woodward, University of Minnesota, report 0.47- to 2.5-μm spectroscopy of V723 Cas (cf. IAUC 8676) on 2007 Dec. 15 UT using VNIRIS on the Lick 3-m telescope. This long-lived, super-soft x-ray source is faint but has retained its high-excitation spectrum, showing emission lines of [Fe X], [Fe XI], and C VI.

M31N 2007-12b Rudy, Lynch, Prater, Mazuk, Gilbert, Puetter, and Perry report that

spectroscopy, obtained as above, of the apparent nova 2007-12b in M31 (cf.

http://www.cfa.harvard.edu/iau/CBAT M31.html) on 2007 Dec. 15 UT

shows a remarkably bright Hα line of very high equivalent width. Owing to the low signal-to-noise ratio on this faint object, no continuum was detectable, and the only other emission lines seen were Paβ, Paγ, Paδ, and He I 1.0830-μm. The FWHM of these lines were about 5000 km/s — notably broad for a nova. The FWZI of Hα was ∼ 6800 km/s.

2008 January 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green